# MarS-FM: Generative Modeling of Molecular Dynamics via Markov State Models

### v2

@@ Abstract
Molecular Dynamics (MD) is a powerful computational microscope for probing protein function. However, the need for fine-grained integration and the long timescales of biomolecular events make MD computationally expensive. To address this, several generative models have been proposed to generate surrogate trajectories at lower cost. Yet, these models typically learn a fixed-lag transition density, causing the training signal to be dominated by frequent but uninformative transitions. We introduce a new class of generative models, MSM Emulators, which instead learn to sample transitions across discrete states defined by an underlying Markov State Model (MSM). We instantiate this class with Markov Space Flow Matching (MarS-FM), whose sampling offers more than two orders of magnitude speedup compared to implicit- or explicit-solvent MD simulations. We benchmark MarS-FM’s ability to reproduce MD statistics through structural observables such as RMSD, radius of gyration, and secondary structure content. Our evaluation spans protein domains (up to 500 residues) with significant chemical and structural diversity, including unfolding events, and enforces strict sequence dissimilarity between training and test sets to assess generalization. Across all metrics, MarS-FM outperforms existing methods, often by a substantial margin.
11footnotetext: Work done as an intern at Valence Labs/Recursion22footnotetext: Corresponding author: kacper.kapusniak@cs.ox.ac.uk
@@ 1 Introduction
Deep Learning has unlocked fast and accurate prediction of proteins’ 3D structures (Jumper et al., 2021; Baek et al., 2021; Krishna et al., 2024; Abramson et al., 2024). However, these methods do not capture dynamic behavior of proteins (Lewis et al., 2024), whose conformational ensembles are governed by the Boltzmann distribution. To study such dynamics, the most reliable computational tool is Molecular Dynamics (MD) (Alder & Wainwright, 1959; Rahman, 1964; McCammon et al., 1977; Risken & Risken, 1996), which simulates atomic motion by integrating Netwon’s Law. Long MD trajectories provide samples from the Boltzmann distribution and reveal the mechanisms of biomolecular interactions, a key asset in drug discovery (De Vivo et al., 2016). Nevertheless, MD is costly, as events in biology occur over timescales vastly longer than the simulation timestep. This challenge has spurred a range of enhanced sampling methods to accelerate dynamics, often through non-physical forces (Laio & Parrinello, 2002; Hamelberg et al., 2004; Jiang & Roux, 2010; Sabbadin & Moro, 2014).
Recent works have proposed to avoid the computational burden of MD by using generative flows to sample from the Boltzmann distribution (Noé et al., 2019; Zheng et al., 2024; Klein & Noe, 2024). A key subclass is the family of MD Emulators (MD-Emus), which learn to emulate MD-sampling by modeling the transition density associated with a fixed lag time $\tau$ (Klein et al., 2023; Schreiner et al., 2023; Nam et al., 2024; Diez et al., 2024; Costa et al., 2024). That is, MD-Emus are models that generate future conformations $x(t+\tau)$ conditioned on some input frame $x(t)$. At inference, these methods are applied autoregressively to produce surrogate MD trajectories. Jing et al. (2024b) advanced this approach by training a model to generate multiple frames jointly, all separated by a fixed interval $\tau$. However, learning dynamics at a fixed lag time introduces key challenges. Short lag times limit the achievable speed-up, while long lag times can skip over important meta-stable states. More fundamentally, the training signal is dominated by frequent but uninformative transitions observed during the simulation, while high-barrier transitions that drive exploration remain underrepresented. As a result, MD-Emus may struggle to capture rare, large conformational changes (Figure 1, left).
Limitations of MD-Emus stem from the data imbalance in MD, whose temporal dynamic contains irrelevant high-frequency information. A common strategy to extract meaningful signal from MD through coarse-grained representations, is offered by Markov State Models (MSMs) (Noé et al., 2009; Prinz et al., 2011; Bowman et al., 2013). MSMs cluster frames into discrete states and describe dynamics via a Markov chain matrix $\mathsf{T}$. While compressing the dynamical content, MSMs guarantee estimation of long-time statistics through the equilibrium distribution induced by $\mathsf{T}$.
@@
Main Contributions. We present MSM Emulators (MSM-Emus), a new class of generative models that learn to sample from the Markov chain transition induced by an MSM, rather than emulating the noisy temporal dynamics induced by MD. This modeling shift has important implications as illustrated in Figure 1. Since target transitions now depend on state connectivity rather than specific observed paths, MSM-Emus benefit from much higher training diversity for learning rare conformational changes across high-energy barriers. In fact, MSM-Emus interpolate across discrete macroscopic states associated with large conformational changes, such as folding or unfolding, rather than among specific frames, which is a more robust and generalizable signal across different proteins. During sampling, frames are more easily decorrelated and autoregressive calls are greatly reduced, mitigating compounding error effects. We showcase this class with Markov Space Flow Matching (MarS-FM), a novel framework optimized using Flow Matching (Lipman et al., 2023).
Evaluation with chemical and structural diversity. MD-Emus have usually been evaluated on small peptides or datasets with limited chemical (Lindorff-Larsen et al., 2011; Majewski et al., 2023) or structural (Vander Meersche et al., 2024) diversity. To address this gap, we evaluate MarS-FM on MD-CATH (Mirarchi et al., 2024), a large-scale dataset of thousands of protein domains with sequence length up to 500 residues. In particular, by leveraging the highest-temperature replica, we are able to test MarS-FM’s ability to capture large conformational changes, such as unfolding. To assess generalization, we adopt the protocol of Lewis et al. (2024) and use mmseqs2 (Steinegger & Söding, 2017) with maximum sensitivity to ensure that test proteins share no more than $20\%$ sequence similarity with any protein in the training set. We benchmark MarS-FM using structural observables reported in Jing et al. (2024b); Mirarchi et al. (2024), including RMSD, radius of gyration, and secondary structure content, and show that it consistently outperforms MD-Emus—often by a large margin. Crucially, MarS-FM explores the target distribution much more efficiently than both MD and MD-Emus, generating conformations across different states even in low-sample regimes.
@@ 2 Preliminaries and Setting
A molecular system of NN atoms, in thermal equilibrium at temperature $T$, explores conformations according to the Boltzmann distribution $\mu$, i.e. $q\sim\mu(q)\propto\exp(-{\mathcal{H}}(q)/(k_{B}T))$ with $q=(x,\xi)\in{\mathcal{X}}$ denoting the system’s position and velocity, $\mathcal{H}$} the Hamiltonian, and $k_{B}$ the Boltzmann constant. We focus on proteins, whose motions—such as local unfolding or cryptic pocket formation—arise from structural fluctuations. To identify these conformational changes and quantify their frequencies (free energy), we aim to sample structures from the Boltzmann distribution $\mu$. This enables the measurement of observables $\phi$—functions defined on the phase space $\mathcal{X}$}—via expectations $\mathbb{E}_{x\sim\mu}[\phi(x)]$.
Since direct sampling from $\mu$ is intractable, a widespread alternative is to use Molecular Dynamics (MD). MD simulates a continuous-time Markov process, designed to be ergodic w.r.t. $\mu$. That is, for any observable, time averages over a long trajectory approximate expectations under $\mu$ (Schütte et al., 2023). A common formulation of MD is via Langevin dynamic: given a system with potential energy $\mathcal{U}$}, position and velocity of an atom $i$ with mass $m_{i}$, are updated as follows
@@
with $\gamma$ and $\sigma$ controlling friction and thermal noise. In practice, we discard the kinetic component and focus on the trajectories of positions $t\mapsto x(t)$. While MD is a powerful tool, its computational cost is a major bottleneck. Many biologically relevant events occur on millisecond timescales, while Langevin dynamics requires integration with time steps of the order of femtoseconds, making long simulations prohibitively expensive. Crucially, meta-stable states of a protein are often separated by high-energy barriers, leading to MD simulations wasting time trapped in deep local minima (Wales, 2005).
@@ 3 MD Emulators and the challenges of fixed lag time transitions
A key object describing MD sampling is the transition density $y\sim p_{\tau}(y|x)$ associated with a lag time $\tau$, which represents the probability of a state $x$ evolving to state $y$ within time $\tau$. Concretely, this is estimated from MD trajectories by examining future transitions $x(t+\tau)$ given an input frame $x(t)$. MD Emulators (MD-Emus) are a class of conditional generative models trained to approximate $p_{\tau}(\cdot|x(t))$ from MD data (Fu et al., 2023; Klein et al., 2023; Schreiner et al., 2023; Nam et al., 2024; Diez et al., 2024; Costa et al., 2024). Specifically, MD-Emus learn to generate future conformations $y\sim p^{\theta}_{\tau}(y|x(t))$ given $x(t)$. Common training approaches include Normalizing Flows (Rezende & Mohamed, 2015; Chen et al., 2018), Flow Matching (Lipman et al., 2022; Tong et al., 2023; Albergo et al., 2023), or Score Matching (Ho et al., 2020; Song et al., 2021).
These models train a neural network $v_{\theta}$conditioned on sequence and an input conformation x(t)x(t)—to match the empirical distribution of transitions observed in the MD data. Jing et al. (2024b) refined such an approach in MDGen, by generating the next $K$ future conformations at once. Explicitly, MDGen learns to sample from the joint density $\mathbf{y}=(y_{1},\ldots y_{K})\sim p_{\tau,K}(\mathbf{y}|x)=\prod_{i}p_{\tau}(y_{i+1}|y_{i})$, conditioned on $y_{0}=x.$ In the following, we assume we are given an MD-Emu that is trained to approximate $p_{\tau,K}$.
@@ 3.1 Limitations of MD-Emus
Training. One of the key challenges in learning from MD trajectories is the data imbalance between frequent but uninformative transitions in an energy minima, and rare but key transitions across different minima (Wales, 2005). To make things more concrete, consider the case where a protein MD trajectory visits two (or more) macroscopic states $S_{A}$ and $S_{B}$, such as folded and unfolded conformations, separated by high-energy barriers. Transitions $S_{A}\rightarrow S_{B}$ are rare but critical as they reveal conformational flexibility and functional dynamics. However in any given trajectory interval of length $K\tau$, the system is far more likely to remain in single state, than to cross an energy barrier. As a result, training batches for MD-Emus are typically dominated by high-frequency, low-information intra-state transitions. This data imbalance limits the model’s ability to learn conditional distributions describing meaningful state changes. For example, to learn how to transition from folded to unfolded states, the model must observe actual transitions $S_{A}\ni x(t)\rightarrow x(t+K\tau)\in S_{B}$ in the data—but such events are rare. This sparsity constrains the diversity and efficiency of training data and affects generalization, as MD-Emus learn to replicate fine-grained temporal resolution rather than large conformational changes.
Inference. MD-Emus need to be used autoregressively to generate long trajectories. This may lead to error accumulation, causing samples to progressively drift away from the data manifold. Nam et al. (2024) introduced a refinement step to address this, yet this comes at the expense of training a second network. Besides, using a larger $K$ does not fully resolve the problem, as the number of required samples often exceeds the maximum window feasible during training for larger proteins.
The challenges faced by MD-Emus stem from fixed lag time transitions often being uninformative and failing to capture meaningful conformational changes across different energy minima. To overcome these limitations, we first need to apply a coarse-grained representation to MD data so to discard high-frequency information. To this aim, we rely on Markov State Models (MSMs) and introduce a novel class of generative flows to learn transitions directly among states (rather than frames).
@@ 4 The new class of MSM Emulators
Markov State Models (MSMs) provide a coarse-grained representation that removes high-frequency time signal while still capturing long-timescale statistics (Noé et al., 2009; Pande et al., 2010; Prinz et al., 2011; Husic & Pande, 2018). In MSMs, frames $x(t)$ are assigned to discrete states $S_{1},\ldots,S_{M}$ and the dynamics is modeled as a Markov chain over these states. Given an interval $\tau$, one estimates a transition matrix $\mathsf{T}$, where $\mathsf{T}_{ij}$ is the probability of transitioning from $S_{i}$ to $S_{j}$ within time $\tau$:
@@
In its simplest form, MSMs assumes a uniform density within each state. That is, given $x(t)\in S_{i}$ and $y\in S_{j}$, $p_{\mathsf{T}}(y|x(t))$ depends only on the identity of the clusters, but not on the specific conformations. This coarse-graining enables robust estimation of long-timescale dynamics, even from limited transition data. In principle, one could initialize short MD simulations (of length $\tau$) across discrete states $\{S_{i}\}$, and propagate their dynamics via $\mathsf{T}$. However, sampling from the states of an underlying (unseen) MSM remains a challenge. We propose a class of generative models that addresses this issue. {defn} MSM Emulators (MSM-Emus) are generative models that are trained to sample conformations from $p_{\mathsf{T}}(\cdot|x(t)))$, where $\mathsf{T}$ is the Markov chain transition matrix of a given MSM. By matching the distribution $p_{\mathsf{T}}(\cdot|x(t))$—that depends only on the MSM states—MSM-Emus bypass many of the data scarcity and imbalance issues that affect MD-Emus §3.
@@
Fast exploration. At inference, MSM-Emus can explore the energy landscape by generating samples $y\sim p_{\mathsf{T}}(y|x(0))$, without being bound by fine-grained temporal dynamics. This significantly improves sampling efficiency (as visualized in Figure 1 and validated in §5). Additionally, as many conformations can now be generated in parallel, MSM-Emus reduce compounding errors due to autoregressive calls.

### v3

@@ Abstract
Molecular Dynamics (MD) is a powerful computational microscope for probing protein function. However, the need for fine-grained integration and the long timescales of biomolecular events make MD computationally expensive. To address this, several generative models have been proposed to generate surrogate trajectories at lower cost. Yet, these models typically learn a fixed-lag transition density, causing the training signal to be dominated by frequent but uninformative transitions. We introduce a new class of generative models, MSM Emulators, which instead learn to sample transitions across discrete states defined by an underlying Markov State Model (MSM). We instantiate this class with Markov Space Flow Matching (MarS-FM), whose sampling offers more than two orders of magnitude speedup compared to implicit- or explicit-solvent MD simulations. We benchmark MarS-FM’s ability to reproduce MD statistics through structural observables such as RMSD, radius of gyration, and secondary structure content. Our evaluation spans protein domains (up to 500 residues) with significant chemical and structural diversity, including unfolding events, and enforces strict sequence dissimilarity between training and test sets to assess generalization. Across all metrics, MarS-FM outperforms existing methods, often by a substantial margin. Code is available at https://github.com/valence-labs/mars-fm.
@@ 1 Introduction
Deep Learning has unlocked fast and accurate prediction of proteins’ 3D structures (jumper2021highly; baek2021accurate; krishna2024generalized; abramson2024accurate). However, these methods do not capture dynamic behavior of proteins (lewis2024scalable), whose conformational ensembles are governed by the Boltzmann distribution. To study such dynamics, the most reliable computational tool is Molecular Dynamics (MD) (alder1959studies; rahman1964correlations; mccammon1977dynamics; risken1996fokker), which simulates atomic motion by integrating Newton’s Law. Long MD trajectories provide samples from the Boltzmann distribution and reveal the mechanisms of biomolecular interactions, a key asset in drug discovery (de2016role). Nevertheless, MD is costly, as events in biology occur over timescales vastly longer than the simulation timestep. This challenge has spurred a range of enhanced sampling methods to accelerate dynamics, often through non-physical forces (laio2002escaping; hamelberg2004accelerated; jiang2010free; sabbadin2014supervised).
Recent works have proposed to avoid the computational burden of MD by using generative flows to sample from the Boltzmann distribution (noe2019boltzmann; zheng2024predicting; kleintransferable). A key subclass is the family of MD Emulators (MD-Emus), which learn to emulate MD-sampling by modeling the transition density associated with a fixed lag time $\tau$ (klein2023timewarp; schreiner2023implicit; nam24generative; diez2024boltzmann; costa2024equijump). That is, MD-Emus are models that generate future conformations $x(t+\tau)$ conditioned on some input frame $x(t)$. At inference, these methods are applied autoregressively to produce surrogate MD trajectories. jing24generative advanced this approach by training a model to generate multiple frames jointly, all separated by a fixed interval $\tau$. However, learning dynamics at a fixed lag time introduces key challenges. Short lag times limit the achievable speed-up, while long lag times can skip over important meta-stable states. More fundamentally, the training signal is dominated by frequent but uninformative transitions observed during the simulation, while high-barrier transitions that drive exploration remain underrepresented. As a result, MD-Emus may struggle to capture rare, large conformational changes (Figure 1, left).
Limitations of MD-Emus stem from the data imbalance in MD, whose temporal dynamic contains irrelevant high-frequency information. A common strategy to extract meaningful signal from MD through coarse-grained representations, is offered by Markov State Models (MSMs) (noe2009constructing; prinz2011markov; bowman2013introduction). MSMs cluster frames into discrete states and describe dynamics via a Markov chain matrix $\mathsf{T}$. While compressing the dynamical content, MSMs guarantee estimation of long-time statistics through the equilibrium distribution induced by $\mathsf{T}$.
@@
Main Contributions. We present MSM Emulators (MSM-Emus), a new class of generative models that learn to sample from the Markov-chain transition induced by an MSM constructed from MD trajectories. In this view, the MSM provides a coarse-grained Markov model of long-timescale dynamics between metastable states, and MSM-Emus focus on learning transitions between metastable states rather than to the full high-frequency MD time series. This modeling shift has important implications as illustrated in Figure 1.
Since target transitions now depend on state connectivity rather than specific observed paths, MSM-Emus benefit from much higher training diversity for learning rare conformational changes across high-energy barriers. In fact, MSM-Emus interpolate across discrete macroscopic states associated with large conformational changes, such as folding or unfolding, rather than among specific frames, which is a more robust and generalizable signal across different proteins. During sampling, frames are more easily decorrelated and autoregressive calls are greatly reduced, mitigating compounding error effects. We showcase this class with Markov Space Flow Matching (MarS-FM), a novel framework optimized using Flow Matching (lipman_flow_2022).
Evaluation with chemical and structural diversity. MD-Emus have usually been evaluated on small peptides or datasets with limited chemical (lindorff2011fast; majewski2023machine) or structural (vander2024atlas) diversity. To address this gap, we evaluate MarS-FM on MD-CATH (mirarchi2024mdcath), a large-scale dataset of thousands of protein domains with sequence length up to 500 residues. In particular, by leveraging the highest-temperature replica, we are able to test MarS-FM’s ability to capture large conformational changes, such as unfolding. To assess generalization, we adopt the protocol of lewis2024scalable and use mmseqs2 (steinegger2017mmseqs2) with maximum sensitivity to ensure that test proteins share no more than $20\%$ sequence similarity with any protein in the training set. We benchmark MarS-FM using structural observables reported in jing24generative; mirarchi2024mdcath, including RMSD, radius of gyration, and secondary structure content, and show that it consistently outperforms MD-Emus—often by a large margin. Crucially, MarS-FM explores the target distribution much more efficiently than both MD and MD-Emus, generating conformations across different states even in low-sample regimes.
@@ 2 Preliminaries and Setting
A molecular system of $N$ atoms, in thermal equilibrium at temperature $T$, explores conformations according to the Boltzmann distribution $\mu$, i.e. $q\sim\mu(q)\propto\exp(-{\mathcal{H}}(q)/(k_{B}T))$ with $q=(x,\xi)\in{\mathcal{X}}$ denoting the system’s position and velocity, $\mathcal{H}$} the Hamiltonian, and $k_{B}$ the Boltzmann constant. We focus on proteins, whose motions—such as local unfolding or cryptic pocket formation—arise from structural fluctuations. To identify these conformational changes and quantify their frequencies (free energy), we aim to sample structures from the Boltzmann distribution $\mu$. This enables the measurement of observables $\phi$—functions defined on the phase space $\mathcal{X}$}—via expectations $\mathbb{E}_{x\sim\mu}[\phi(x)]$.
Since direct sampling from $\mu$ is intractable, a widespread alternative is to use Molecular Dynamics (MD). MD simulates a continuous-time Markov process, designed to be ergodic w.r.t. $\mu$. That is, for any observable, time averages over a long trajectory approximate expectations under $\mu$ (schutte2023overcoming). A common formulation of MD is via Langevin dynamic: given a system with potential energy $\mathcal{U}$}, position and velocity of an atom $i$ with mass $m_{i}$, are updated as follows
@@
with $\gamma$ and $\sigma$ controlling friction and thermal noise. In practice, we discard the kinetic component and focus on the trajectories of positions $t\mapsto x(t)$. While MD is a powerful tool, its computational cost is a major bottleneck. Many biologically relevant events occur on millisecond timescales, while Langevin dynamics requires integration with time steps of the order of femtoseconds, making long simulations prohibitively expensive. Crucially, meta-stable states of a protein are often separated by high-energy barriers, leading to MD simulations wasting time trapped in deep local minima (wales2005energy).
@@ 3 MD Emulators and the challenges of fixed lag time transitions
A key object describing MD sampling is the transition density $y\sim p_{\tau}(y|x)$ associated with a lag time $\tau$, which represents the probability of a state $x$ evolving to state $y$ within time $\tau$. Concretely, this is estimated from MD trajectories by examining future transitions $x(t+\tau)$ given an input frame $x(t)$. MD Emulators (MD-Emus) are a class of conditional generative models trained to approximate $p_{\tau}(\cdot|x(t))$ from MD data (fusimulate; klein2023timewarp; schreiner2023implicit; nam24generative; diez2024boltzmann; costa2024equijump). Specifically, MD-Emus learn to generate future conformations $y\sim p^{\theta}_{\tau}(y|x(t))$ given $x(t)$. Common training approaches include Normalizing Flows (rezende2015variational; chen2018neural), Flow Matching (lipman_flow_2022; tong2023improving; albergo2023stochastic), or Score Matching (ho2020denoising; song2020score).
These models train a neural network $v_{\theta}$conditioned on sequence and an input conformation x(t)x(t)—to match the empirical distribution of transitions observed in the MD data. jing24generative refined such an approach in MDGen, by generating the next $K$ future conformations at once. Explicitly, MDGen learns to sample from the joint density $\mathbf{y}=(y_{1},\ldots y_{K})\sim p_{\tau,K}(\mathbf{y}|x)=\prod_{i}p_{\tau}(y_{i+1}|y_{i})$, conditioned on $y_{0}=x.$ In the following, we assume we are given an MD-Emu that is trained to approximate $p_{\tau,K}$.
@@ 3.1 Limitations of MD-Emus
Training. One of the key challenges in learning from MD trajectories is the data imbalance between frequent but uninformative transitions in an energy minima, and rare but key transitions across different minima (wales2005energy). To make things more concrete, consider the case where a protein MD trajectory visits two (or more) macroscopic states $S_{A}$ and $S_{B}$, such as folded and unfolded conformations, separated by high-energy barriers. Transitions $S_{A}\rightarrow S_{B}$ are rare but critical as they reveal conformational flexibility and functional dynamics. However in any given trajectory interval of length $K\tau$, the system is far more likely to remain in single state, than to cross an energy barrier. As a result, training batches for MD-Emus are typically dominated by high-frequency, low-information intra-state transitions. This data imbalance limits the model’s ability to learn conditional distributions describing meaningful state changes. For example, to learn how to transition from folded to unfolded states, the model must observe actual transitions $S_{A}\ni x(t)\rightarrow x(t+K\tau)\in S_{B}$ in the data—but such events are rare. This sparsity constrains the diversity and efficiency of training data and affects generalization, as MD-Emus learn to replicate fine-grained temporal resolution rather than large conformational changes.
Inference. MD-Emus need to be used autoregressively to generate long trajectories. This may lead to error accumulation, causing samples to progressively drift away from the data manifold. nam24generative introduced a refinement step to address this, yet this comes at the expense of training a second network. Besides, using a larger $K$ does not fully resolve the problem, as the number of required samples often exceeds the maximum window feasible during training for larger proteins.
The challenges faced by MD-Emus stem from fixed lag time transitions often being uninformative and failing to capture meaningful conformational changes across different energy minima. To overcome these limitations, we first need to apply a coarse-grained representation to MD data so as to discard high-frequency information. To this aim, we rely on Markov State Models (MSMs) and introduce a novel class of generative flows to learn transitions directly among states (rather than frames). The key advantage is that, once trajectories have been clustered into metastable states, we are no longer restricted to training only on frame pairs of the form $x_{t}\rightarrow x_{t+\tau}$ obtained by slicing the time series at a fixed resolution. Instead, any pair of frames assigned to successive MSM states can be used for training, greatly increasing the number and diversity of informative transitions.
@@ 4 The new class of MSM Emulators
Markov State Models (MSMs) provide a coarse-grained representation that removes high-frequency time signal while still capturing long-timescale statistics (noe2009constructing; pande2010everything; prinz2011markov; husic2018markov). In MSMs, frames $x(t)$ are assigned to discrete states $S_{1},\ldots,S_{M}$ and the dynamics is modeled as a Markov chain over these states. Given an interval $\tau$, one estimates a transition matrix $\mathsf{T}$, where $\mathsf{T}_{ij}$ is the probability of transitioning from $S_{i}$ to $S_{j}$ within time $\tau$:
@@
In its simplest form, MSMs assumes a uniform density within each state. That is, given $x(t)\in S_{i}$ and $y\in S_{j}$, $p_{\mathsf{T}}(y|x(t))$ depends only on the identity of the clusters, but not on the specific conformations. This coarse-graining enables robust estimation of long-timescale dynamics, even from limited transition data. In principle, one could initialize short MD simulations (of length $\tau$) across discrete states $\{S_{i}\}$, and propagate their dynamics via $\mathsf{T}$. However, sampling from the states of an underlying (unseen) MSM remains a challenge. We propose a class of generative models that addresses this issue.
Definition 1.
MSM Emulators (MSM-Emus) are generative models that are trained to sample conformations from $p_{\mathsf{T}}(\cdot|x(t)))$, where $\mathsf{T}$ is the Markov chain transition matrix of a given MSM.
For a frame $x(t)\in S_{i}$, the MSM defines a categorical distribution over successor states $j\mapsto{\mathsf{T}}_{ij}$. We interpret $p_{\mathsf{T}}(\cdot\mid x(t))$ as the resulting mixture distribution over MD frames: one first samples a destination state $S_{j}$ according to $\mathsf{T}_{ij}$ and then samples a conformation from the empirical ensemble of MD frames assigned to $S_{j}$. MSM-Emus are thus trained to draw conformations from this MSM-induced mixture, rather than to learn the exact frame-to-frame MD transition density. By matching the distribution $p_{\mathsf{T}}(\cdot|x(t))$—that depends only on the MSM states—MSM-Emus bypass many of the data scarcity and imbalance issues that affect MD-Emus §3.
@@
Fast exploration. At inference, MSM-Emus can explore the energy landscape by generating samples $y\sim p_{\mathsf{T}}(y|x(0))$, without being bound by fine-grained temporal dynamics. This significantly improves sampling efficiency (as visualized in Figure 1 and validated in §5). Additionally, as many conformations can now be generated in parallel, MSM-Emus reduce compounding errors due to autoregressive calls. While MSM-Emus are designed to preserve the coarse-grained kinetics encoded by the MSM (e.g., metastable-state populations and transition probabilities), they are not intended to faithfully reproduce all fine-grained dynamical observables of the MD trajectories, at the same level of accuracy as fixed-lag MD-Emus. By construction, training on state-to-state jump transitions prioritizes long-timescale behavior over the detailed intra-state time ordering. In this sense, MSM-Emus are complementary to MD-Emus: the former emphasize efficient exploration and long-timescale thermodynamics/kinetics, whereas the latter more directly track the local, frame-to-frame dynamics (see Appendix C.2 for a concrete hybrid instantiation and torsion-angle decorrelation analysis).